### v1

# Stochastic resonance in vibrational polariton chemistry

Abstract
In this work, we systematically investigate the impact of ambient noise intensity on the rate modifications of ground-state chemical reactions in an optical cavity under vibrational strong-coupling conditions. To achieve this, we utilize a numerically exact open quantum system approach–the hierarchical equations of motion in twin space, combined with a flexible tree tensor network state solver. Our findings reveal a stochastic resonance phenomenon in cavity-modified chemical reactivities: an optimal reaction rate enhancement occurs at an intermediate noise level. In other words, this enhancement diminishes if ambient noise, sensed by the cavity-molecule system through cavity leakage, is either too weak or excessively strong. In the collective coupling regime, when the cavity is weakly damped, rate enhancement strengthens as more molecules couple to the cavity. In contrast, under strong cavity damping, reaction rates decline as the number of molecules grows.
preprint: AIP/123-QED
I Introduction
Placing reactive molecules within an optical microcavity under conditions of vibrational strong coupling (VSC) has recently shown promise for manipulating chemical reactions, such as cavity catalysis and cavity-induced reaction selectivities, even in the absence of light.Ebbesen (2016) This concept has been demonstrated in several experiments, which suggests that adjusting the distance between two reflective dielectric mirrors in a Fabry-Pérot cavity so that the cavity frequency resonates with specific molecular vibrations can alter reaction rates or product ratios.Thomas et al. (2016); Vergauwe et al. (2019); Lather et al. (2019); Thomas et al. (2019); Hiura, Shalabney, and George (2019); Thomas et al. (2020); Hirai, Hutchison, and Uji-i (2020); Hirai et al. (2020); Sau et al. (2021); Lather et al. (2022); Ahn et al. (2023); Ebbesen et al. (2023) However, two independent experimental attempts to replicate these findings, while successfully reproducing VSC conditions, did not observe noticeable rate changes in an on-resonant cavity.Imperatore, Asbury, and Giebink (2021); Wiesehan and Xiong (2021) Different outcomes in the experiments suggest that additional factors beyond VSC, factors that may have been previously overlooked, could play a critical role in achieving cavity-induced modifications of ground-state chemical reactivities. In this study, we aim to explore one such experimental variable: the damping strength of the cavity mode. This damping, caused by ambient noise surrounding the cavity due to the unavoidable cavity leakage, may influence the extent of cavity-induced rate modifications.
Naturally, this intriguing yet controversial phenomenon has spurred a substantial amount of theoretical research.Galego, Garcia-Vidal, and Feist (2016, 2017); Galego et al. (2019); Campos-Gonzalez-Angulo and Yuen-Zhou (2020); Mandal, Montillo Vega, and Huo (2020); Yang and Cao (2021); Li, Mandal, and Huo (2021a); Li, Nitzan, and Subotnik (2020); Li, Mandal, and Huo (2021b); Sun and Vendrell (2022); Schäfer et al. (2022); Mandal, Li, and Huo (2022); Lindoy, Mandal, and Reichman (2022); Wang et al. (2022); Fischer, Anders, and Saalfrank (2022); Fiechter et al. (2023); Sokolovskii and Groenhof (2023); Pavošević, Smith, and Rubio (2023) Recent studies using quantum dynamical simulations underscore the necessity of a quantum discrete state description for both molecular vibrations and the cavity mode.Lindoy, Mandal, and Reichman (2023, 2024); Ying and Huo (2023); Hu, Ying, and Huo (2023); Ke, Borrelli, and Thoss (2022); Ke and Richardson (2024) However, this requirement introduces significant challenges, particularly in the collective regime, where a large ensemble of molecules couples to the cavity mode. To address the computational challenges, we employ the hierarchical equations (HEOM) of motion along with a tree tensor network state (TTNS) solver.Ke (2023) The HEOM method is a well-established, numerically exact approach for open quantum system dynamics,Tanimura and Kubo (1989); Yan et al. (2004); Ishizaki and Tanimura (2005); Xu and Yan (2007); Shi et al. (2009); Yan (2014); Jin, Zheng, and Yan (2008); Schinabeck, Härtle, and Thoss (2018); Hsieh and Cao (2018); Shi et al. (2018); Tanimura (2020) enabling a non-perturbative and non-Markovian treatment of system dynamics and bath-related observables. The TTNS,Shi, Duan, and Vidal (2006); Tagliacozzo, Evenbly, and Vidal (2009); Murg et al. (2010); Li, von Delft, and Xiang (2012); Changlani et al. (2013); Nakatani and Chan (2013); Lubich et al. (2013); Murg et al. (2015); Gunst et al. (2018); Schröder et al. (2019); Larsson (2019); Ferrari, Magnifico, and Montangero (2022); Montangero, Rico, and Silvi (2022); Sulz et al. (2024) in addition, provides an efficient data compression scheme for storing and propagating the high-dimensional composite system-plus-bath wavefunction.
To gain a preliminary understanding of how chemical reactions inside a cavity respond to ambient noise–stemming from interactions between the confined cavity mode and the continuum of far-field electromagnetic modes, which dampens the cavity’s oscillatory dynamics–we begin our study on the single-molecule level. Our results show that reaction rate enhancements within the cavity exhibit a typical stochastic resonance featureGammaitoni et al. (1998): neither excessively weak nor overly strong damping of cavity dynamics supports effective cavity-induced rate modifications. Instead, the rate constant inside a resonant cavity reaches its peak at an intermediate level of cavity damping. In the collective coupling regime, where multiple molecules are interconnected through a shared coupling with the cavity mode, rate modifications can display distinct behaviors as the system size grows. Under weak ambient noise, the rate enhancement strengthens as more molecules are coupled to the cavity mode. On the contrary, under conditions of strong cavity damping, the rate enhancement is attenuated with the increased number of coupled molecules. These observations lead us to ask an essential question: Could an ensemble of molecules on the macroscopic scale, immersed in a solvent and thus exposed to significant background noise, act cooperatively through collective coupling to the cavity mode, and synergetically harness a feeble external fluctuation to optimize the reaction efficiency?
The following sections present a detailed account of our findings. Sec. II outlines an open quantum system model that describes chemical reactions in a condensed-phase optical cavity and the quantum dynamics methodology used. Sec. III presents and discusses the numerical results, followed by a summary and perspectives for our future research in Sec. IV.
II Theory
We consider an open quantum system model to study chemical reactions inside an optical cavity,
$H=H_{\rm S}+H_{\rm E},$ (1)
where $H_{\rm S}$ describes the cavity-molecule system and $H_{\rm E}$ represents its surrounding environment.
The reaction dynamics of $N_{\rm mol}$ molecules coupled to a single-mode cavity is described by the Pauli-Fierz light-matter Hamiltonian in the dipole gauge under the long-wavelength approximation Flick et al. (2017); Rokaj et al. (2018); Mandal et al. (2023); Lindoy, Mandal, and Reichman (2023)
$\begin{split}H_{\rm S}=&\overbrace{\frac{p_{\rm c}^{2}}{2}+\frac{1}{2}\omega_{\rm c}^{2}\left(x_{\rm c}+\sqrt{\frac{2}{\omega_{\rm c}}}\eta_{\rm c}\sum_{i=1}^{N_{\mathrm{mol}}}\vec{\mu}_{i}(x_{i})\cdot\vec{e}\right)^{2}}^{H_{\rm c}(\mathbf{x})}\\ &+\sum_{i=1}^{N_{\rm mol}}\underbrace{\left[\frac{p_{i}^{2}}{2}+\frac{E_{\mathrm{b}}}{a^{4}}(x_{i}-a)^{2}(x_{i}+a)^{2}\right]}_{H_{\rm mol}^{i}}.\end{split}$ (2)
The cavity mode is modeled as a harmonic oscillator with momentum $p_{\rm c}$, coordinate $q_{\rm c}$, and frequency $\omega_{\rm c}$. For convenience, we set $\hbar=1$ throughout this work. The light-matter coupling strength is characterized by the parameter $\eta_{\rm c}=\frac{1}{\omega_{\rm c}}\sqrt{\frac{\omega_{\rm c}}{2\epsilon_{0}V}}$, where $\epsilon_{0}$ is the permittivity of the medium within the cavity and $V$ is the quantization volume of the electromagnetic mode. The unit vector $\vec{e}$ points to the light polarization direction. Each molecule is represented by a reactive vibrational mode with mass-scaled momentum $p_{i}$, coordinate $x_{i}$, and a symmetric double-well potential energy surface with a barrier height $E_{\rm b}$ between two local minima at $x_{i}=\pm a$. The molecular dipole moment, denoted by $\vec{\mu}_{i}(x_{i})$, depends on the reaction coordinate $x_{i}$.
The solvent and the continuum of electromagnetic modes together constitute the dissipative environment, with the Hamiltonian given by
$\begin{split}H_{\rm E}=&\sum_{k}\frac{P_{{\rm c}k}^{2}}{2}+\frac{1}{2}\omega_{{\rm c}k}^{2}\left(Q_{{\rm c}k}+\frac{g_{{\rm c}k}}{\omega_{{\rm c}k}^{2}}x_{\rm c}\right)^{2}\\ &+\sum_{i=1}^{N_{\rm mol}}\sum_{k}\frac{P^{2}_{ik}}{2}+\frac{1}{2}\omega_{ik}^{2}\left(Q_{ik}+\frac{g_{ik}x_{i}}{\omega_{ik}^{2}}\right)^{2},\end{split}$ (3)
where the cavity mode and each molecule are coupled to their respective baths, represented as collections of harmonic oscillators. The bath associated with the cavity mode is referred to as the cavity bath, and each molecular bath is referred to as a solvent bath. The $k$th oscillator in bath $\alpha$ (either cavity or solvent) is described by its momentum $P_{\alpha k}$ and coordinate $Q_{\alpha k}$, with frequency $\omega_{\alpha k}$ and coupling strength $g_{\alpha k}$. The coupling to the system induces a displacement $g_{\alpha k}x_{\alpha}/\omega_{\alpha k}^{2}$ in each oscillator’s coordinate.
The environmental influence on the system dynamics is fully captured by the time-correlation function
$C_{\alpha}(t)=\frac{1}{\pi}\int_{-\infty}^{\infty}\frac{e^{-i\omega t}}{1-e^{-\beta\omega}}J_{\alpha}(\omega)\mathrm{d}\omega=\sum_{p=1}^{P\rightarrow\infty}\lambda_{\alpha}\eta_{\alpha p}e^{-i\gamma_{\alpha p}t},$ (4)
where $\beta=1/k_{\rm B}T$ is the inverse temperature, and $J_{\alpha}(\omega)=\frac{\pi}{2}\sum_{k}\frac{c_{\alpha k}^{2}}{\omega_{\alpha k}}\delta(\omega-\omega_{\alpha k})$ is the spectral density function of bath $\alpha$. This spectral density encodes the coupling-weighted density of states in the frequency domain for each bath. The time correlation function $C_{\alpha}(t)$ can be expanded analytically or numerically as a series of exponentials.Hu, Xu, and Yan (2010); Xu et al. (2022) This exponential expansion can be understood as mapping the original bath onto an effective one, which can be truncated efficiently at finite temperatures. In this effective representation, the bath consists of $P$ dissipative bosonic modes, each with a complex-valued frequency $\gamma_{\alpha p}$ and a component-wise coupling strength $\eta_{\alpha p}$ to the system. The reorganization energy $\lambda_{\alpha}=\sum_{k}\frac{g_{\alpha k}^{2}}{2\omega_{\alpha k}^{2}}$ quantifies the overall coupling strength between the system and bath $\alpha$.
Rooted in the HEOM method in twin space for the system,Borrelli (2019); Borrelli and Gelin (2021); Ke, Borrelli, and Thoss (2022) we can formulate a Schrödinger equation as
$\frac{\mathrm{d}|\Psi(t)\rangle}{\mathrm{d}t}=-i\mathcal{H}|\Psi(t)\rangle$ (5)
for the extended wavefunction of the composite system
$|\Psi(t)\rangle\equiv\sum_{\mathbf{n}}\sum_{v_{\rm c}v^{\prime}_{\rm c}\cdots v_{N_{\rm mol}}v^{\prime}_{N_{\rm mol}}}C^{\mathbf{n}}_{v_{\rm c}v^{\prime}_{\rm c}\cdots v_{N_{\rm mol}}v^{\prime}_{N_{\rm mol}}}(t)|v_{\rm c}v^{\prime}_{\rm c}\cdots v_{N_{\rm mol}}v^{\prime}_{N_{\rm mol}}\rangle\otimes|\mathbf{n}\rangle.$ (6)
which includes both the system (cavity and molecules) and the effective baths. Each physical DoF in the system is represented by two independent indices, $v_{j}$ and $v^{\prime}_{j}$, which arise from the purification of the density matrices. A single operator $O_{j}$ for the $j$th system DoF in Hilbert space corresponds to a pair of superoperators in twin space: $\hat{O}_{j}=O_{j}\otimes I_{j}$ and $\tilde{O}_{j}=I_{j}\otimes O_{j}^{\dagger}$, where $I_{j}$ is the identity operator for the $j$th DoF. Further, $|\mathbf{n}\rangle=|\cdots n_{\alpha p}\cdots\rangle$ represents the Fock state for the effective environmental modes in the number representation, with creation and annihilation operators $b_{\alpha p}^{+}$ and $b_{\alpha p}$ defined as
$b_{\alpha p}^{+}|\mathbf{n}\rangle=\sqrt{n_{\alpha p}+1}|\mathbf{n}_{\alpha p}^{+}\rangle;$ (7a)
$b_{\alpha p}|\mathbf{n}\rangle=\sqrt{n_{\alpha p}}|\mathbf{n}_{\alpha p}^{-}\rangle,$ (7b)
where $|\mathbf{n}_{\alpha p}^{\pm}\rangle=|\cdots n_{\alpha p}\pm 1\cdots\rangle$. The super-Hamiltonian $\mathcal{H}$ is non-Hermitian and can be explicitly expressed as
$\begin{split}\mathcal{H}=&\hat{H}_{\rm S}+\sum_{\alpha}\lambda_{\alpha}\hat{x}_{\alpha}^{2}-\tilde{H}_{\rm S}-\sum_{\alpha}\lambda_{\alpha}\tilde{x}_{\alpha}^{2}-i\sum_{\alpha}\sum_{p}\gamma_{\alpha p}b_{\alpha p}^{+}b_{\alpha p}\\ &+\sum_{\alpha}\sum_{p}\sqrt{\lambda_{\alpha}}\left[(\hat{x}_{\alpha}-\tilde{x}_{\alpha})b_{\alpha p}+(\eta_{\alpha p}\hat{x}_{\alpha}-\eta^{*}_{\alpha p}\tilde{x}_{\alpha})b_{\alpha p}^{+}\right].\end{split}$ (8)
By projecting Eq. (5) onto a specific environmental Fock state, we obtain a differential equation for the auxiliary density operator, $\dot{\rho}^{\mathbf{n}}(t)=\langle\mathbf{n}|\dot{\Psi}(t)\rangle$, which reproduces the standard form of the HEOM method (see a review paper in Ref. Tanimura, 2020 and the literature therein). In particular, the reduced system density operator is given by $\rho_{\rm S}(t)=\rho^{\mathbf{0}}(t)=\langle\mathbf{n}=\mathbf{0}|\Psi(t)\rangle$. In this work, we assume that the system and the environment are initially factorized, so that $\rho(0)=\rho_{\rm S}(0)\cdot\rho_{\rm E}(0)$. Additionally, the environment is assumed to be in thermal equilibrium at the moment $t=0$. This assumption is equivalent to setting the environmental Fock state to the ground state at the initial time in the HEOM formalism.
a) Refer to caption
b) Refer to caption
c) Refer to caption
d) Refer to caption
Figure 1: a) Schematic illustration of an open quantum system model for studying chemical reactions within an optical cavity. The system comprises a single-mode cavity field interacting with $N_{\rm mol}$ molecules. In the illustrated example, $N_{\rm mol}=8$. Each molecule is surrounded by a multitude of solvent molecules, while the confined cavity mode is also influenced by far-field radiation modes outside the cavity. b) and c) are graphical representations of a TTNS decomposition of the extended wavefunction $|\Psi(t)\rangle$ and $|{\rm II}\rangle$, respectively. Each colored circular node in the TTNS represents a low-rank tensor associated with a specific component of the model: molecules (in red), solvent (in green), cavity mode (in yellow), and cavity bath (in blue). Each molecular node is linked to the cavity node through three connecting nodes shown in gray. d) represents a TTNO decomposition of the super-Hamiltonian $\mathcal{H}$. The square nodes represent the low-rank tensors in the TTNO, and they share the same color-coding scheme as in the TTNS. However, unlike in the TTNS representation, physical nodes in the TTNO have two dangling legs instead of one. The numbers labeled next to the connecting edges specify the bond dimensions of the virtual indices. The parameters $d_{\rm m}$, $d_{\rm c}$, and $d_{\rm e}$ indicate the number of basis functions for the vibrational, cavity, and effective environmental DoFs.
To facilitate efficient propagation of the Schrödinger equation in Eq. (5), we employ a tree tensor network state solver.Ke (2023) The high-rank coefficient tensor $C^{\mathbf{n}}_{v_{\rm c}v^{\prime}_{\rm c},\cdots v_{N_{\rm mol}}v^{\prime}_{N_{\rm mol}}}(t)$ is decomposed as a tree tensor network state, which is a powerful way to handle the complicated tensor structure involved in this system. As an example, a graphical representation of the TTNS decomposition of $|\Psi(t)$ for a system with $N_{\rm mol}=8$ molecules and $P=4$ effective environmental modes per bath is shown in Fig. 1(b). In this decomposition, a node with $z$-legs corresponds to a rank-$z$ tensor. To achieve an optimal balance between numerical stability, computational efficiency, and simulation time, the number of legs $z$ is constrained to $z\leq 3$. The physical indices $v_{\rm c}/v^{\prime}_{\rm c}$ (for the cavity), $v_{i}/v^{\prime}_{i}$ (for the molecules), and $n_{{\rm c}p},\,n_{ip}$ (for the effective environmental modes) in $|\Psi(t)\rangle$ are represented by the dangling legs in the yellow (filled/hollow), red (filled/hollow), blue, and green nodes, respectively. The connected legs between neighboring nodes $j$ and $k$ correspond to virtual indices $r_{jk}$, which runs from $1$ to $D_{jk}$. The maximum value among the bond dimensions $\{D_{jk}\}$ is called the maximal bond dimension $D_{\rm max}$. In practice, $D_{\rm max}$ is systematically increased until the results converge to a desired accuracy. The gray nodes, which have exclusively connected legs, are termed connecting nodes. The root node is assigned to the filled yellow node for the cavity, which has no parent node. All other nodes, except for the root node, have a parent node. This parent node is the adjacent node on the path pointing toward the root. In the binary tree structure displayed in Fig. 1(b), each node has at most two children. The leaves are the nodes without any offspring, representing the termination points of the tree.
The super-Hamiltonian $\mathcal{H}$ can be represented as a tree tensor network operator (TTNO) with the same tree topology as the TTNS used for $|\Psi(t)\rangle$. This decomposition ensures that when the TTNO is applied to the TTNS of $\Psi(t)$, the tree structure is preserved. The construction of the TTNO for $\mathcal{H}$ can be performed in an automatic manner.Milbradt, Huang, and Mendl (2024a, b); Ceruti, Kressner, and Sulz (2024); Li et al. (2024) In this work, we use an optimal TTNO constructor based on bipartite theory, which allows for an efficient construction of the TTNO.Li et al. (2024) The resulting TTNO, corresponding to the tree shape shown in Fig. 1(b), is depicted in Fig. 1(d). A key difference between the TTNS and the TTNO is that in the TTNO, the physical nodes possess two dangling legs. To evolve Eq. (5), we adopt a time propagation scheme based on the time-dependent variational principle (TDVP), as detailed in Refs. Haegeman et al., 2016; Paeckel et al., 2019; Dunnett and Chin, 2021; Bauernfeind and Aichhorn, 2020; Kloss, Reichman, and Bar Lev, 2020; Ceruti, Lubich, and Walach, 2021. After solving Eq. (5), observables of interest can be extracted from the wavefunction $|\Psi(t)\rangle$.
For the symmetric double-well model considered in this work, we assume that reactant and product regions are divided by a surface at $x_{i}^{\ddagger}=0$. The population of the $i$th molecule in the product region is obtained as
$P_{i}(t)=\mathrm{Tr}\left\{h_{i}\rho(t)\right\}=\langle{\rm II}|\hat{h}_{i}|\Psi(t)\rangle.$ (9)
Here $h_{i}=\theta(x_{i}-x_{i}^{\ddagger})$ is the Heaviside projection operator onto the product region, where $x_{i}>x_{i}^{\ddagger}$. The wavefunction $|{\rm II}\rangle=|1_{\mathrm{sys}}\rangle\otimes|\mathbf{n=0}\rangle$ corresponds to the state where the system is in a unit vector state in twin space, i.e. , $|1_{\mathrm{sys}}\rangle=\otimes_{i}\sum_{v_{i}=v^{\prime}_{i}}|v_{i}v^{\prime}_{i}\rangle$, and the environmental Fock state is in the ground state.
To perform the inner product in Eq. (9), the wavefunction $|{\rm II}\rangle$ is also decomposed as a TTNS, maintaining the same tree topology as $|\Psi(t)\rangle$, as displayed in Fig. 1(c). The number next to the legs denotes the bond dimension of the corresponding index. The parameters $d_{\rm m}$, $d_{\rm c}$, and $d_{\rm e}$ represent the number of basis functions for the molecular reactive DoF, cavity mode, and the effective environmental bosons, respectively. Each node in the TTNS is represented by a rank-$3$ tensor $V_{i_{1}\times i_{2}\times i_{3}}$, where $i_{1}$, $i_{2}$, and $i_{3}$ are the dimensions of the respective indices. For the connecting nodes, the tensor is $V_{1\times 1\times 1}=1$. For the molecular (in red) or cavity (in yellow) nodes, the non-zero entries in the tensor are $V_{1\times d_{\rm m/c}\times d_{\rm m/c}}[1,j,k]=\delta_{jk}$ (filled node) and $V_{d_{\rm m/c}\times 1\times d_{\rm m/c}}[j,1,k]=\delta_{jk}$ (hollow node). For the environmental nodes (in green and blue), the non-zero entry is $V_{1\times 1\times d_{\rm e}}[1,1,1]=1$.
The rigorous expression for the first-order reaction rate constant of the $i$th molecule, moving from the reactant (left well) to the product region (right well), in the flux correlation function formalism is given byMiller, Schwartz, and Tromp (1983); Craig, Thoss, and Wang (2007); Chen and Shi (2009); Ke et al. (2022)
$k_{i}=\lim_{t\rightarrow t_{p}}k_{i}(t)=\lim_{t\rightarrow t_{p}}\frac{C_{i}^{\mathrm{fs}}(t)}{1-2P_{i}(t)}.$ (10)
Here, $t_{p}$ is the time at which $k_{i}(t)$ reaches a plateau. The flux-side correlation function is defined as
$C_{i}^{\mathrm{fs}}(t)=\mathrm{Tr}\{\rho(t)F_{i}\}=\langle{\rm II}|\tilde{F}_{i}|\Psi(t)\rangle,$ (11)
where $F_{i}=i[H,h_{i}]$ is the flux operator.
In the context of quantum dynamical simulations, it is important to note that the plateau value of $k_{i}(t)$ in Eq. (10), which corresponds to the reaction rate, is independent of the choice of the initial density matrix. For the calculations in this work, the initial molecular vibrational wavefunction is largely prepared in the reactant region, and the cavity mode is assumed to be in thermal equilibrium. The system density operator is then defined as
$\rho_{\rm S}(0)=\left(\prod_{i=1}^{N_{\rm mol}}\frac{e^{-\frac{\beta H^{i}_{\rm mol}}{2}}(1-h_{i})e^{-\frac{\beta H^{i}_{\rm mol}}{2}}}{Z_{i}}\right)\cdot\frac{e^{-\beta H_{\rm c}(\mathbf{x}=0)}}{\mathrm{Tr}\left\{e^{-\beta H_{\rm c}(\mathbf{x}=0)}\right\}}$ (12)
with $Z_{i}=\mathrm{Tr}_{i}\left\{e^{-\frac{\beta H^{i}_{\rm mol}}{2}}(1-h_{i})e^{-\frac{\beta H^{i}_{\rm mol}}{2}}\right\}$ is a partition function for the $i$th molecule. The singular value decomposition of the initial density matrix in Eq. (12) is employed to construct the initial tree tensor network state.
In the simulations conducted for this study, the following parameters are used for the double-well model, $E_{b}=2250\,\text{cm}^{-1}$ and $a=44.4$ a.u. , which are consistent with previous studies in the field.Lindoy, Mandal, and Reichman (2023, 2024); Ying and Huo (2023); Hu, Ying, and Huo (2023); Ke, Borrelli, and Thoss (2022); Ke and Richardson (2024); Fiechter et al. (2023) The Debye-Lorentzian spectral density function is adopted to characterize the baths and their interaction with the system
$J_{\alpha}(\omega)=\frac{2\lambda_{\alpha}\omega\Omega_{\alpha}}{\omega^{2}+\Omega_{\alpha}^{2}},$ (13)
[…]
It is worth noting that while the matrix product state (MPS),Schollwöck (2011) also known as tensor train (TT),Oseledets (2011) is a particular, simpler case of the TTNS, using MPS/TT for simulating reaction dynamics in vibrational polariton chemistry–especially in the collective regime–is computationally inefficient. This inefficiency arises due to the complexity of correlations in such systems, which require more flexible tensor network topologies. Further details on the numerical performance of different tensor network structures are provided in Appendix A.
a) $\eta_{c}=0.00125$ a.u.
b) $\eta_{c}=0.005$ a.u.
Figure 2: Ratio of rates $k_{\rm ic}/k_{\rm oc}$ for a single molecule undergoing a reaction inside an optical cavity $(k_{\rm ic})$ compared to the outside rate $(k_{\rm oc})$ as a function of the cavity damping strength $\lambda_{\rm c}$. The results for two different temperatures $T$, characteristic frequencies $\Omega_{\rm c}$ of the cavity bath, and light-matter coupling strengths ($\eta_{\rm c}=0.00125$ a.u. in the upper panel and $\eta_{\rm c}=0.005$ a.u. in the lower panel) are shown. The cavity frequency is kept at $\omega_{\rm c}=1185\,\text{cm}^{-1}$ and $d_{\rm c}=10$.
III Results
In practice, the cavity is not perfectly reflective and is therefore subject to external noise from the continuum of far-field electromagnetic modes. Previous theoretical studies have demonstrated that, in a resonant but lossless cavity, reaction rates remain unaffected.Lindoy, Mandal, and Reichman (2023); Ying and Huo (2023); Ke, Borrelli, and Thoss (2022) This finding gives a hint that cavity leakage may be a crucial factor in altering reaction dynamics inside an optical microcavity. Accordingly, the objective of this work is to understand how ground-state chemical reactions within an optical cavity are influenced by cavity damping. We begin by investigating the effect of cavity damping strength on reaction rate modifications at the single-molecule level, then extend our analysis to the collective regime, where multiple molecules are simultaneously coupled to the cavity mode.
In this work, cavity leakage is modeled by introducing a cavity bath, with the cavity damping strength quantified by the parameter $\lambda_{\rm c}$. This is verified by examining the dynamics of the average photonic excitation number, given by $\langle n_{\rm c}(t)\rangle=\mathrm{Tr}\{a^{\dagger}_{\rm c}a_{\rm c}\rho(t)\}$. Here, $a^{\dagger}_{\rm c}=\sqrt{\omega_{\rm c}/2}(x_{\rm c}-ip_{\rm c/\omega_{\rm c}})$ and $a_{\rm c}=\sqrt{\omega_{\rm c}/2}(x_{\rm c}+ip_{\rm c/\omega_{\rm c}})$ represent the creation and annihilation operators of the confined cavity mode, respectively.
a) Refer to caption
b) Refer to caption
Figure 3: a) Ratio of rates $k_{\rm ic}/k_{\rm oc}$ as a function of the number of molecules collectively coupled to the cavity mode. Note that $k_{\rm ic}/k_{\rm oc}$ is identical for each individual molecule as the constituent of the molecular ensemble. Results are shown for different numbers of $d_{\rm c}$ lowest photonic states. The cavity frequency is fixed at $\omega_{c}=1185\,\text{cm}^{-1}$. b) Rate modification profile as a function of the cavity frequency for cases with one and four molecules coupled to the cavity mode, respectively, and with $d_{\rm c}=6$. Remaining parameters are $\lambda_{\rm c}=50\,\text{cm}^{-1}$, $\Omega_{\rm c}=1000\,\text{cm}^{-1}$, $\eta_{\rm c}=0.00125$ a.u. , and $T=300\,\text{K}$.
a) Refer to caption
b) Refer to caption
Figure 4: Same as Fig. 3 expect that $\lambda_{\rm c}=165\,\text{cm}^{-1}$.
Fig. 2 presents the cavity-induced rate modification, represented as $k_{\rm ic}/k_{\rm oc}$–the ratio of reaction rates inside and outside the cavity–in the single-molecule limit, plotted as a function of $\lambda_{\rm c}$. The cavity frequency is set at $\omega_{\rm c}=1185\,\text{cm}^{-1}$, closely resonant with molecular vibrational transitions and corresponding to the peak in the rate modification profile over $\omega_{\rm c}$ (see also Fig. 3b and Fig. 4b). The rate within the cavity initially increases with $\lambda_{\rm c}$, reaching a maximum before this rate enhancement diminishes as $\lambda_{\rm c}$ continues to increase. This behavior is examined with various values of temperature $T$, the characteristic frequency $\Omega_{\rm c}$ of the cavity bath, and the light-matter coupling strength $\eta_{\rm c}$. The extent of the rate modification depends on temperature, $\Omega_{\rm c}$, and $\eta_{\rm c}$. Notably, the cavity-induced rate modification is a finite-temperature effect, vanishing at low temperatures and saturating as temperature rises, as demonstrated in our previous study.Ke, Borrelli, and Thoss (2022) Furthermore, the rate modification increases with $\Omega_{\rm c}$ until it reaches a saturation point. However, both temperature $T$ and $\Omega_{\rm c}$ only slightly affect the crossover value of $\lambda_{\rm c}$. In contrast, increasing the light-matter coupling strength $\eta_{\rm c}$ not only strengthens the rate modification significantly but also shifts the crossover $\lambda_{\rm c}$ to a much higher value, rendering the rate modification more resilient to ambient noise.
This observation fits into the notion of the stochastic resonance theory, which was first proposed to explain the periodic recurrence of Earth’s ice ages.Benzi, Sutera, and Vulpiani (1981); Nicolis (1981) Over time, it has prospered into a broad field on its own with diverse applications across physics, chemistry, biomedical sciences, engineering, and beyond, highlighting the constructive role of an optimal amount of noise.Wiesenfeld and Moss (1995); Gammaitoni et al. (1998); Andò and Graziani (2000); Hänggi (2002); Harmer, Davis, and Abbott (2002); Wellens, Shatokhin, and Buchleitner (2003); Moss, Ward, and Sannita (2004) Stochastic resonance refers to a phenomenon where a moderate level of noise improves a system’s sensitivity to detect weak, information-carrying signals.Gammaitoni et al. (1998) For chemical reactions inside an optical cavity, three key elements have been possessed for the emergence of stochastic resonance. Firstly, reactions take place in a solvent, which serves as a source of substantial background noise. Secondly, reactive molecules are inherently nonlinear systems that support threshold-like barrier-crossing events. Thirdly, cavity leakage exposes reactions to a generically weak coherent input, which could boost the system’s sensitivity.
In many cases, stochastic resonance is observed as a collective effect.Jung et al. (1992); Lindner et al. (1995) For example, it is found that the sensitivity of paddlefish electroreceptors to detect individual plankton is enhanced when plankton form a swarm.Russell, Wilkens, and Moss (1999) Similarly, it is of great interest to explore how the reaction dynamics inside the cavity change in the collective regime, by gradually increasing the number of molecules coupled to the cavity mode.
We begin with a weakly-damped cavity mode characterized by $\lambda_{\rm c}=50\,\text{cm}^{-1}$, with other parameters set as $T=300\,\text{K}$ and $\Omega_{\rm c}=1000\,\text{cm}^{-1}$. In addition, the light-matter coupling strength per molecule is fixed at $\eta_{\rm c}=0.00125$ a.u. , meaning that the Rabi splitting $\Omega_{R}$ increases as the number of molecules grows. For a homogeneous molecular aggregate, the rate ratio $k_{\rm ic}/k_{\rm oc}$ is identical for every individual molecule in the ensemble. In Fig. 3a, we show $k_{\rm ic}/k_{\rm oc}$ as a function of the number of molecules collectively coupled to the cavity mode, with $d_{\rm c}$ lowest photonic states included in the calculations. As more molecules are coupled through the cavity mode, the reaction rate of an individual molecule as part of the connected molecular network inside the cavity increases. It is worth noting that, in the collective regime, the involvement of higher photonic excitation manifolds becomes necessary for accurate results. For the scenario in Fig. 3, at least the two lowest photonic excited states $(d_{\rm c}=3)$ are required to reach convergence. If only the lowest photonic excited state $(d_{\rm c}=2)$ is considered, both the reaction rates and the decay of oscillations in the flux-side correlation functions and $\langle n_{\rm c}(t)\rangle$ are significantly underestimated, suggesting the importance of multi-photon processes in capturing the dissipative nature of the cavity bath, which is crucial for cavity-induced reaction modifications. Furthermore, we observed that as the number of molecules increases from $N_{\rm mol}=1$ to $N_{\rm mol}=4$, the height of the resonant peak in the reaction rate modification profile increases, as shown in Fig. 3b, while the peak becomes sharper. Specifically, the full width at half maximum (FWHM) is reduced from approximately $125\,\text{cm}^{-1}$ for $N_{\rm mol}=1$ to $105\,\text{cm}^{-1}$ for $N_{\rm mol}=4$.
In contrast, when the cavity mode is more heavily damped with $\lambda_{\rm c}=165\,\text{cm}^{-1}$ while all other parameters remain unchanged, increasing $N_{\rm mol}$ results in a steady drop of $k_{\rm ic}/k_{\rm oc}$, as shown in Fig. 4a. Additionally, in this scenario, the FWHM of the resonant peak in the cavity-induced rate modification profile broadens from about $80\,\text{cm}^{-1}$ for $N_{\rm mol}=1$ to $105\,\text{cm}^{-1}$ for $N_{\rm mol}=4$, as shown in Fig. 4b. Moreover, with the larger $\lambda_{\rm c}$, more photonic excited states contribute to the collective reaction dynamics within an optical cavity. For instance, $d_{\rm c}=4$ photonic states are required for $N_{\rm mol}>4$, as illustrated in Fig. 4a.
These findings support the hypothesis that cavity-induced rate modifications might also be influenced by factors beyond just the presence of VSC, typically indicated by a large Rabi splitting. It is known that the Rabi splitting $\Omega_{R}$ is proportional to the molecular dipole moment $\vec{\mu}_{i}(x_{i})$ and the square root of molecular concentration $\sqrt{N_{\rm mol}/V}$, while remaining largely independent of other experimental factors, such as broadening line width and cavity quality factor $Q$.Ebbesen (2016) In our single-molecule limit results, the Rabi splitting is held constant. However, the reaction rate is considerably impacted by the cavity damping strength, which is highly associated with the specifics of the cavity setup, including the material composition and quality factor. In the collective coupling scenario, where the Rabi splitting increases with the number of molecules coupled to the cavity mode, the rate changes exhibit different trends depending on the cavity damping strength. This suggests that the Rabi splitting alone is not the determining factor for observing rate modification within the cavity. Therefore, our findings may provide an important clue into discrepancies in experimental outcomes.Hiura, Shalabney, and George (2019); Lather et al. (2019); Imperatore, Asbury, and Giebink (2021); Wiesehan and Xiong (2021)
We also notice that in the collective coupling regime, while the reaction rate varies with $N_{\rm mol}$, it does not exceed the maximum rate observed at the single-molecule level under the same conditions of temperature, $\Omega_{\rm c}$, and $\eta_{\rm c}$. This suggests that studies at the single-molecule level may provide valuable predictions on the upper limit of cavity-induced rate modification for a specific chemical reaction in the collective coupling regime.
From another perspective, our results clearly show that the reaction rate for a single molecule, as the building block of a molecular network, can differ significantly from that of an isolated molecule. Moreover, the reaction kinetics of a molecular assembly organized by the cavity mode, in response to a certain level of external noise, are highly sensitive to the system’s size. This raises a compelling question: For a macroscopically large ensemble of molecules collectively coupled to a cavity mode, could even a very weak perturbation surrounding the cavity be perceived and amplified throughout the entire connected cavity-molecule system, leading to a drastic change in reaction dynamics? Validating this conjecture will require further investigations, both theoretically and experimentally.
IV Conclusion
In summary, we conduct a quantum dynamical and numerically exact study on the impact of cavity damping strength on condensed-phase chemical reactions inside an optical microcavity under VSC conditions, using the HEOM method with an efficient tree tensor network decomposition scheme. Our results reveal that cavity-induced rate modifications exhibit a typical stochastic resonance feature. Specifically, reaction rates within a cavity are significantly enhanced when the cavity frequency is tuned in resonance with the molecular vibrational transition energy, not only under VSC conditions but also at an intermediate level of cavity damping. Additionally, distinct behaviors are observed as the number of molecules collectively coupled to the cavity mode increases, depending on the cavity damping strength: under weak damping, reaction rates inside the cavity can be boosted as the number of molecules grows, whereas strong damping leads to a decrease in reaction rates with molecular aggregation.
We hope this study will inspire deeper curiosity into the intricacies of vibrational polariton chemistry experiments, and into factors beyond VSC that could potentially lead to major ramifications in the reaction dynamics. Looking forward, we plan to develop a more advanced tensor network state approach that extends beyond the tree tensor network topology, allowing us to investigate more realistic multi-mode cavities. We also aim to extend this approach to realistic molecular systems by combining it with the ab initio simulations, moving beyond simplified models to more accurately represent experimental conditions.
a)
b) MPS ($1^{\rm st}$ index ordering)
c) MPS ($2^{\rm nd}$ index ordering)
d) TTNS Refer to caption
Figure 5: a) Illustration of MPS/TT decompositions with two different arrangements of physical indices on the left side, and a TTNS decomposition on the right side of the extended wavefunction $|\Psi(t)\rangle$ for $N_{\rm mol}=4$. b-d) Rescaled flux-side correlation function $k_{1}(t)$ for the first molecule, as defined in Eq. (10), is shown with varying maximal bond dimensions. Each panel corresponds to a tensor network state decomposition scheme illustrated in a). Results in c) for $D_{\rm max}=20$, $D_{\rm max}=30$, and $D_{\rm max}=40$ are represented by solid, dashed, and dotted lines, respectively, to avoid visual overlap from superimposition. The pertinent parameters are set as follows: $\omega_{\rm c}=1185\,\text{cm}^{-1}$, $\lambda_{\rm c}=165\,\text{cm}^{-1}$, $\Omega_{\rm c}=1000\,\text{cm}^{-1}$, $\eta_{\rm c}=0.00125$ a.u. , $d_{\rm c}=6$, and $T=300\,\text{K}$.
Acknowledgements.
The author appreciates valuable discussions with Richard Milbradt on the algorithm for constructing a TTNO with state diagrams, and with Prof. Jiushu Shao on the stochastic resonance theory. Y.K. thanks the Swiss National Science Foundation for the award of a research fellowship.
Data Availability Statement
The data and code that support the findings of this work are available from the corresponding author upon reasonable request.
Appendix A Comparison of numerical performances between MPS and TTNS
In this appendix, we demonstrate the advantages of using the TTNS approach over the MPS/TT for studying chemical reactions inside an optical cavity. MPS/TT represents a special, simplified form of TTNS, where all nodes are arranged in a one-dimensional chain, as illustrated in Fig. 5. Although MPS/TT restricts the network topology, its numerical performance is still heavily influenced by the ordering of physical indices. Fig. 5b and Fig. 5c compare $k(t)$ obtained with two different index orderings in the MPS. In the first configuration, molecular nodes and their associated solvent bath nodes are attached subsequently to the cavity and cavity bath nodes on the right. While each molecule couples equally to the cavity mode, the higher-index molecular nodes are positioned farther from the cavity node, giving rise to long-distance entanglement and thus necessitating a considerably large bond dimension to accurately capture the system dynamics. This issue can be partially mitigated by arranging half of the molecular and associated bath nodes on either side of the cavity nodes, as shown in Fig. 5c. Nevertheless, a large bond dimension is still required to obtain smooth and converged results.
In the TTNS decomposition scheme tailored to the collective regime of vibrational polariton chemistry, each molecular node is positioned equidistant from the cavity node, separated by only a few connecting nodes. This tree topology and index arrangement significantly reduce the need for large bond dimensions, optimizing computational efficiency while preserving accuracy. For instance, simulations with a bond dimension of $D_{\rm max}=20$ are sufficient to produce high-quality results, as shown in Fig. 5d.